# Explainable Graph Neural Networks Under Fire

### v1

@@ 5 Experimental Results and Analysis
Before answering the research questions, we state three presumptions and check them empirically in Section 5.1. Next, we answer RQ1 and RQ2 in Section 5.2, and RQ3 in Section 5.3.
5.1 Experiments to Check Presumptions
@@
Presumption 1. The prediction correctness of a GNN predictor and explanation accuracy of a GNN explainer are not related. For each dataset, we first get the set of correctly predicted nodes and the set of wrongly predicted nodes; then we compare the distribution of their explanation accuracy. As shown in Figure 2 in Appendix E, there is no notable difference. In other words, we do not observe that wrongly labelled nodes tend to have lower explanation accuracy despite Faber et al. (2021) claim that it is unfair to use wrongly labelled nodes to evaluate GNN explanations (as they may not use the ground-truth explanations to make predictions and thus tend to have lower explanation accuracy).
@@
Presumption 2. The prediction confidence of a GNN predictor and explanation accuracy of a GNN explainer are related, where prediction confidence is defined as the probability assigned to the predicted class. For each dataset, we first divide the nodes into five groups based on their prediction confidence; then we compare the distribution of their explanation accuracy. Tables 410 in Appendix E show that group ‘G9’ (i.e., the set of nodes with prediction confidence in $[0.9,1.0]$) indeed tends to have lower explanation accuracy on the original graph than other groups.
@@
Presumption 3. The prediction confidence of a GNN predictor and attack performance of a GNN explanation attacker are related. To check this presumption, we first divide the nodes into five groups based on their prediction confidence; then we compare the distribution of their attack performance. Tables 410 in Appendix E show that group ‘G9’ also tends to have less successful attacks (i.e., smaller $\Delta\text{GEA}$) when compared to other groups. Possible reasons are that: 1) the nodes in this group have lower explanation accuracy on original graph; and 2) the prediction confidence is close to 1, leaving no much room for perturbations. These two facts together make the attack more challenging.
5.2 Experiments to Answer RQ1 and RQ2: Attack Effectiveness
@@
Table 1: Results on all synthetic datasets (average $\pm$ standard deviations across 5 trials). O. GEA means explanation accuracy on original graph and P. GEA indicates explanation accuracy on perturbed graph. Moreover, $\Delta$GEA is the explanation accuracy change after perturbation, Simcos is the cosine similarity between explanations (before and after perturbations), and #Pert. represents the number of flipped edges. $\Delta$Label is the average change of predicted labels, while $\Delta$Prob denotes the average absolute change of predicted probability (of the original predicted class). “GXAttack", “Rnd. Flipping", and “Rnd. Rewiring" correspond to our attack method, random flipping, and random rewiring, respectively. $\uparrow^{\prime}$ indicates larger values are preferred while $\downarrow^{\prime}$ means the opposite.
@@
From Table 1, we have the following main observations:
@@
Observation 2. The type of motifs (namely ground-truth explanations) appears to have minimal impact on both explanation accuracy and attack performance. This can be obtained by comparing the results on Syn3 (Motif =“House") with those on Syn4 (Motif =“Circle").
@@
Answer to RQ1. From Table 1, we can see that GXAttack can achieve a $\Delta\text{GEA}$ ranging from $0.088$ to $0.267$ while using a very small perturbation budget (less than 7.1 on all datasets) and keeping the predicted labels nearly unchanged (i.e., $\Delta\text{Label}$ is nearly zero) after perturbation.
@@
5.3 Experiments to Answer RQ3: Attack Transferability
@@
Answer to RQ3. From Table 2, we can see that some post-hoc explainers, including Grad, GuidedBP, PGMExplainer, and IG, suffer from poor explanation accuracy on the original graphs, even giving lower accuracy than the Random Explainer. As a result, investigating the attack transferability on these explainers may not be meaningful. On the contrary, like PGExplainer, explainers such as GradCAM, GNNExplainer, and SubgraphX (gray) can achieve good performance on most original graphs. For those, the attacks optimized for PGExplainer are also harmful, leading to a dramatic degradation of explanation accuracy. For instance, the explanation accuracy of GNNexplainer decreases from 0.711 to 0.397 on Syn2. This demonstrates the transferability of the attacks generated by GXAttack.
Other Analysis. We performed a sensitivity analysis for GXAttack in terms of maximal attack budget, maximal training epochs, and the loss trade-off hyperparameter. The results and analysis are deferred to Appendix I due to space reasons. Moreover, we also conducted some property analysis of GXAttack in terms of the prediction confidence, and node degree in Appendix H.
@@
Table 2: Transferability of attacks on selected synthetic datasets with a typical run. The results are consistent on other datasets and thus are deferred to Table 12 to save space. We report the explanation accuracy (i.e., GEA) on the original graph $\rightarrow$ explanation accuracy on the perturbed graph.
Explainer Syn1 Syn2 Syn3 Syn4
PGExplainer 0.728 $\rightarrow$ 0.510 0.648 $\rightarrow$ 0.347 0.485 $\rightarrow$ 0.304 0.471 $\rightarrow$ 0.311
Grad 0.025 $\rightarrow$ 0.065 $0$rightarrow 0.001 $0$rightarrow 0.000 0.003 $\rightarrow$ 0.001
GradCAM 0.745 $\rightarrow$ 0.551 0.708 $\rightarrow$ 0.397 0.510 $\rightarrow$ 0.300 0.511 $\rightarrow$ 0.310
GuidedBP $0$rightarrow 0.000 $0$rightarrow 0.000 $0$rightarrow 0.000 $0$rightarrow 0.000
IG 0.118 $\rightarrow$ 0.104 0.134 $\rightarrow$ 0.103 0.136 $\rightarrow$ 0.101 0.109 $\rightarrow$ 0.092
GNNExplainer 0.738 $\rightarrow$ 0.550 0.711 $\rightarrow$ 0.397 0.539 $\rightarrow$ 0.341 0.514 $\rightarrow$ 0.327
SubgraphX 0.113 $\rightarrow$ 0.015 0.532 $\rightarrow$ 0.375 0.296 $\rightarrow$ 0.141 0.417 $\rightarrow$ 0.282
PGMExplainer 0.015 $\rightarrow$ 0.064 0.046 $\rightarrow$ 0.039 0.031 $\rightarrow$ 0.034 0.036 $\rightarrow$ 0.041
RandomExplainer 0.184 $\rightarrow$ 0.146 0.170 $\rightarrow$ 0.096 0.135 $\rightarrow$ 0.102 0.124 $\rightarrow$ 0.090
6 Related Work
We only briefly revisit closely related work due to space constraints. See Appendix F for more detail.
GNN Explanations. Yuan et al. (2022); Li et al. (2022); Kakkad et al. (2023) perform comprehensive surveys and propose excellent taxonomies of GNN explanation methods. Specifically, GNN explanations methods mainly include: 1) Gradients-based methods: Grad (Simonyan et al., 2013), GradCAM (Pope et al., 2019), GuidedBP (Baldassarre and Azizpour, 2019), Integrated Gradients (IG) (Sundararajan et al., 2017); 2) Decomposition-based methods: GraphLRP (Schwarzenberg et al., 2019), GNN-LRP (Schnake et al., 2021); 3) Surrogate-based methods: GraphLIME (Huang et al., 2022), PGM-Explainer (Vu and Thai, 2020); 4) Generation-based methods: XGNN (Yuan et al., 2020); 5) Perturbation-based methods: GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), GraphMask (Schlichtkrull et al., 2020), SubgraphX (Yuan et al., 2021); and 6) CF-based methods: CF-GNNExplainer (Lucic et al., 2022), RCExplainer (Bajaj et al., 2021).
Robustness/Stability of GNN Explanations. Yuan et al. (2022); Agarwal et al. (2022) point out that a reliable GNN explainer should exhibit stability, namely minor perturbations that do not impact the model predictions should not largely change the explanations. Particularly, Agarwal et al. (2022) theoretically analyze the stability of three GNN explainers, including vanilla Grad, GraphMask and GraphLIME, for which they derive upper bounds on the instability of explanations. Moreover, Bajaj et al. (2021); Fang et al. (2024); Wang et al. (2024); Fan et al. (2023) are closely related but are different from our work (as compared in Section 1 and detailed in Appendix F).
Adversarial Robustness of GNNs. Starting with the seminal works (Zügner et al., 2018; Dai et al., 2018), that both proposed adversarial attacks on GNNs, a rich literature formed in the realm of GNN-specific adversarial attacks, defenses, and certifications (Jin et al., 2021; Günnemann, 2022). Most attacks primarily focus on altering predictions via perturbations of the discrete graph structure (edge insertions/deletions), either between existing nodes (Zügner et al., 2018) or via the insertion of new (adversarial) nodes (Zou et al., 2021). Optimization-based attacks either operate globally (number of edge insertions/deletions) (Xu et al., 2019; Geisler et al., 2021), alter a local neighborhood (Zügner et al., 2018), or can handle a combination of both constraints (Gosch et al., 2023). While adversarial attacks on GNN explanation methods may benefit from optimization methods over the discrete graph structure, a careful discussion and derivation of appropriate attack objectives and metrics was missing.
7 Conclusions
@@
The limitations and broader impact of our work are discussed in Appendices J and  K, respectively.
@@ Acknowledgements
Zhong Li and Matthijs van Leeuwen: this publication is part of the project Digital Twin with project number P18-03 of the research programme TTW Perspective, which is (partly) financed by the Dutch Research Council (NWO). Simon Geisler: This research was supported by the Helmholtz Association under the joint research school “Munich School for Data Science – MUDS".
Appendix for paper submission Explainable Graph Neural Networks Under Fire
Appendix A Revisiting PGExplainer
For completeness, we revisit PGExplainer as follows. For brevity, we denote $\mathcal{G}(v)$ as $\mathcal{G}$ (and $\mathcal{G}_{s}(v)$ as $\mathcal{G}_{s}$ analogously) without causing confusions throughout this subsection.
Formalisation. Given a graph $\mathcal{G}$ and the target node $v$, to explain why $f_{\theta}(\mathcal{G}(v))=c$, PGExplainer will provide a local explanation (i.e. explanatory subgraph) $\mathcal{G}_{s}(v)=(\mathbf{A}_{s},\mathbf{X}_{s})$, where $\mathbf{A}_{s}$ is the subgraph structure and $\mathbf{X}_{s}$ is the node features associated with nodes residing in this subgraph. Specifically, PGExplainer maximises the mutual information between the GNN classifier’s prediction $\mathbf{y}$ and the explanatory subgraph $\mathcal{G}_{s}$:
$\underset{\mathcal{G}_{s}}{\max}~{}\text{MI}(\mathbf{y},\mathcal{G}_{s}):=H(\mathbf{y})-H(\mathbf{y}|\mathcal{G}_{s})$ (17)
where $H(\mathbf{y})$ is the entropy of $f_{\theta}(\mathcal{G})$, and $H(\mathbf{y}|\mathcal{G}_{s})$ is the conditional entropy of the GNN’s prediction on the explanatory graph, namely $f_{\theta}(\mathcal{G}_{s})$. As $\theta$ and $\mathcal{G}$ are fixed at the explanation stage (i.e. $\mathbf{y}$ is also fixed), maximising Eq 17 is equivalent to minimising $H(\mathbf{y}|\mathcal{G}_{s})$.
Relaxations. However, directly minimising $H(\mathbf{y}|\mathcal{G}_{s})$ is intractable as there are $\mathcal{O}(2^{|\mathcal{E}|})$ candidates for $\mathcal{G}_{s}$. Therefore, they considered three relaxations: 1) they assume that the explanatory subgraph $\mathcal{G}_{s}$ is a Gilbert random graph, namely the selections of edges from $\mathcal{G}$ are conditionally independent to each other. As a result, the probability of graph $\mathcal{G}$ can be factorised as $P(\mathcal{G})=\prod_{(i,j)\in\mathcal{E}}P(e_{ij})$. Moreover, they instantiate $P(e_{ij})$ using a Bernoulli distribution $B(\lambda_{ij})$, namely $e_{ij}\sim B(\lambda_{ij})$ with $P(e_{ij}=1)=\lambda_{ij}$. As a result, Eq 17 is equivalent to:
$\underset{\mathcal{G}_{s}}{\min}~{}H(\mathbf{y}|\mathcal{G}_{s})=\underset{\mathcal{G}_{s}}{\min}~{}\mathbb{E}_{\mathcal{G}_{s}}[H(\mathbf{y}|\mathcal{G}_{s})]\approx\underset{\Lambda}{\min}~{}\mathbb{E}_{\mathcal{G}_{s}\sim q(\Lambda)}[H(\mathbf{y}|\mathcal{G}_{s})],$ (18)
where $q(\Lambda)$ describes the distribution of explanatory subgraph $\mathcal{G}_{s}$ that is parameterized by $\lambda$’s in $\Lambda$. 2) they relax the edge weights from binary variables $\{0,1\}$ to continuous variable $[0,1]$ using reparameterization trick. As a result, they can utilise gradient-based methods to optimize the objective function. More specifically, the continuous edge weight $\hat{e}_{ij}$ is calculated as $\hat{e}_{ij}=sigmoid[(\ln(\epsilon)-\ln(1-\epsilon)+\omega_{ij})/\tau]$, where $\epsilon\sim\text{U}(0,1)$, $\tau$ is the temperature hyperparameter, and $\omega_{ij}$ is the learnable parameter in $\Omega$. Hence, Eq 18 is equivalent to:
$\underset{\Omega}{\min}~{}\mathbb{E}_{\epsilon\sim\text{U}(0,1)}[H(\mathbf{y}|\mathcal{G}_{s})].$ (19)
3) for efficiency, they replace conditional entropy $H(\mathbf{y}|\mathcal{G}_{s})$ with cross entropy $H(\mathbf{y},\mathbf{y}_{s})$, and further apply Monte Carlo to approximate it:
$\underset{\Omega}{\min}~{}\frac{-1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\mathcal{G}_{s}^{(k)})),$ (20)
where $K$ is the number of sampled explanatory subgraph $\mathcal{G}_{s}^{(k)}$, and $\theta$ is the set of parameters in trained $f_{\theta}(\cdot)$.
Generalisations. If we consider $\Omega$ in Eq 20 as independent variables, then we have to retrain it each time when we want to explain an unseen instance. To mitigate this issue, they employ a parameterized network to learn $\Omega$, namely $\Omega=h_{\Phi}(\mathcal{G},\mathbf{Z})$, where $\mathbf{Z}$ is the learned node embedding matrix before inputting to the downstream tasks (namely the result of component 1 in GNN classifier). Besides, $h(\cdot)$ is the explanation network that is shared by all edges, and $\Phi$ is the set of its parameters. As a result, in the collective setting Eq 20 becomes:
$\underset{\Phi}{\min}~{}\sum_{i\in\mathcal{I}}\sum_{k=1}^{K}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G}^{(i)})\ln(P_{\theta}(y=c|\mathcal{G}_{s}^{(i,k)})),$ (21)
where $\mathcal{G}^{(i)}$ is the $i$-th input graph from a set $\mathcal{I}$, and $\mathcal{G}_{s}^{(i,k)}$ is its $k$-th sampled explanatory subgraph.
Finally, they implement the explanatory network $h(\cdot)$ as a MLP, which returns the edge importance as explanations. For node classification, the explanation is $\omega_{ij}=\text{MLP}_{\Phi}([\mathbf{z}_{i},\mathbf{z}_{j};\mathbf{z}_{v}])$ for node $v$.
Appendix B Projected Gradient Descent
@@
After each gradient update $\tilde{\mathbf{P}}_{t}\in\mathbb{R}^{n\times n}$ although it must be in $[0,1]^{n\times n}$ for its interpretation as Bernoulli random variable. Moreover, we must ensure that sampling $\mathbf{P}_{t}\in\{0,1\}^{n\times n}\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{t})$ will yield a discrete perturbation $\hat{\mathbf{A}}=\mathbf{A}\oplus\mathbf{P}_{t}$ obeying the admissible perturbations $\|\mathbf{A}-\hat{\mathbf{A}}\|_{0}<B$ with sufficient likelihood. Following Xu et al. (2019); Geisler et al. (2021), we ensure that $\mathbb{E}_{\mathbf{P}\sim\text{Bernoulli}(\tilde{\mathbf{P}_{t}})}[\|\mathbf{P}\|_{0}]\leq B$ and then obtain the final perturbation form a small set of samples. To guarantee that $\tilde{\mathbf{P}}_{t}$ parametrizes a valid distribution and that we obey the budget in expectation, we employ the projection $\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$.
@@
$\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}})\leq B]}(\tilde{\mathbf{P}})=\arg\min_{\tilde{\mathbf{S}}}\|\tilde{\mathbf{P}}-\tilde{\mathbf{S}}\|_{F}^{2}\qquad\text{subject to }\;$ $\sum_{i,j}\tilde{S}_{i,j}\leq B$ (22)
$\;\tilde{\mathbf{S}}\in[0,1]^{n\times n}$
@@
for clipping to the hyperplane and projecting back towards the simplex using a bisection search. See Xu et al. (2019) for details and the derivation.
Appendix C Complexity Analysis of the GXAttack Algorithm
Algorithmic Components
  1. 1.
    Initial Predictions and Explanations:
    • $f_{\theta}(\mathbf{A},\mathbf{X})$: Predicting using the GNN on the graph $\mathcal{G}$. The complexity depends on the GNN architecture but typically involves matrix multiplications between the adjacency matrix $\mathbf{A}$ and the feature matrix $\mathbf{X}$.
    • $g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X}))$: Generating explanations based on the GNN predictions. The complexity is generally proportional to that of the GNN prediction if the explainer works by perturbing inputs and observing changes in output.
  2. 2.
    Loop for Gradient Updates:
    • Prediction and Explanation on Perturbed Graph: This operation is similar in complexity to the initial prediction and explanation but is repeated $T$ times (number of epochs).
    • Gradient Computation and Update:
      • Calculating the gradient $\nabla_{\tilde{\mathbf{P}}_{t-1}}\mathcal{L}$ involves computing derivatives of the loss function with respect to the perturbation matrix $\tilde{\mathbf{P}}$. This can be expensive as it depends on the sensitivity of both the GNN predictions and explanations to changes in the adjacency matrix.
      • Updating $\tilde{\mathbf{P}}_{t}$ involves a simple addition and a projection operation. The projection step is constrained by the perturbation budget $B$, but the actual operation is typically linear with respect to the number of potential edges $n^{2}$.
  3. 3.
    Binary Perturbation Sampling:
    • Sampling from a Bernoulli distribution for each potential edge (pair of nodes), which has a complexity proportional to $n^{2}$.
Computational Complexity
We discuss the dense case here and refer to the summary below for a discussion of GNNs/explainers implemented with sparse matrices.
  1. 1.
  2. 2.
    Each Epoch:
    • Prediction and explanation on perturbed graph: $O(2T\cdot(f_{\theta}+g_{\lambda}))$.
  3. 3.
    Sampling Step: $O(n^{2})$.
Space Complexity
  1. 1.
  2. 2.
Summary
Assuming GNN and explainer are in $O(f_{\theta})=O(g_{\lambda})=O(Ed+nd^{2})$ with number of edges $E$, a dense attack would dominate the overall complexity ($O(n^{2})$). An immediate remedy would yield PRBCD (Geisler et al., 2021) since its complexity is $O(B+E)$ with edge flipping budget $B$ and all additional components (e.g., loss function) could be implemented in $O(E)$ with sparse matrices. This yields an overall time and space complexity of $O(E)$, assuming $B\ll E$ and hidden dimensions $d\ll n$, which is the same asymptotic complexity as the sparse GNN $f_{\theta}$ and explainer $g_{\lambda}$.
Appendix D Datasets Description
Overall, we consider the following synthetic datasets, which are summarised in Table 3: 1) synthetic datasets with different size via varying the following hyper-parameters in ShapeGGen: num_subgraphs for the number of subgraphs and subgraph_size for the number of expected nodes in each subgraph; 2) synthetic datasets with different types of explanatory graphs: shape as “circle" or “house"; and 3) synthetic datasets with different node degrees: prob_connection with different values.
Table 3: Summary of synthetic datasets with ground-truth explanations. #Nodes and #Edges represent the number of expected nodes and edges, respectively. We do not consider larger datasets for two reasons: 1) explanation accuracy will be very low ($<0.3$) for most explainers, and it is not meaningful to attack low accuracy explainers; 2) computation will be slow for the attacker (and the explainer).
Dataset Syn1 Syn2 Syn3 Syn4 Syn5 Syn6 Syn7
Shape House House House Circle House House House
#Subgraphs 10 50 50 50 50 50 100
Subgraph Size 6 6 10 10 10 10 10
P(Connection) 0.3 0.06 0.06 0.06 0.12 0.20 0.06
#Classes 2 2 2 2 2 2 2
#Nodes 59 299 521 511 489 492 1018
#Edges 164 970 1432 1314 1664 1898 3374
Average Degree 2.8 3.2 2.7 2.6 3.4 3.9 3.3
Max Budget 10 15 15 15 15 15 15
Appendix E Presumptions Checking Results

### v2

@@ VI Experimental Results and Analysis
Before answering the research questions, we state three presumptions and check them empirically in Section VI-A. Next, we answer RQ1 and RQ2 in Section VI-B, and RQ3 in Section VI-C. Following this, we perform sensitivity analysis in Section VI-D and property analysis in Section VI-E.
VI-A Experiments to Check Presumptions
@@
Presumption 1. The prediction correctness of a GNN predictor and explanation accuracy of a GNN explainer are not related. For each dataset, we first get the set of correctly predicted nodes and the set of wrongly predicted nodes; then we compare the distribution of their explanation accuracy. As shown in Figure 6 in Appendix, there is no notable difference. In other words, we do not observe that wrongly labelled nodes tend to have lower explanation accuracy despite [49] claim that it is unfair to use wrongly labelled nodes to evaluate GNN explanations (as they may not use the ground-truth explanations to make predictions and thus tend to have lower explanation accuracy).
@@
Presumption 2. The prediction confidence of a GNN predictor and explanation accuracy of a GNN explainer are related, where prediction confidence is defined as the probability assigned to the predicted class. For each dataset, we first divide the nodes into five groups based on their prediction confidence; then we compare the distribution of their explanation accuracy. Table IV in Appendix show that group ‘G9’ (i.e., the set of nodes with prediction confidence in $[0.9,1.0]$) indeed tends to have lower explanation accuracy on the original graph than other groups.
@@
Presumption 3. The prediction confidence of a GNN predictor and attack performance of a GNN explanation attacker are related. To check this presumption, we first divide the nodes into five groups based on their prediction confidence; then we compare the distribution of their attack performance. Table IV in Appendix show that group ‘G9’ also tends to have less successful attacks (i.e., smaller $\Delta\text{GEA}$) when compared to other groups. Possible reasons are that: 1) the nodes in this group have lower explanation accuracy on original graph; and 2) the prediction confidence is close to 1, leaving no much room for perturbations. These two facts together make the attack more challenging.
VI-B Experiments to Answer RQ1 and RQ2: Attack Effectiveness
@@
TABLE II: Results on all synthetic datasets (average $\pm$ standard deviations across 5 trials). O. GEA means explanation accuracy on original graph and P. GEA indicates explanation accuracy on perturbed graph. Moreover, $\Delta$GEA is the explanation accuracy change after perturbation, Simcos is the cosine similarity between explanations (before and after perturbations), and #Pert. represents the number of flipped edges. $\Delta$Label is the average change of predicted labels, while $\Delta$Prob denotes the average absolute change of predicted probability (of the original predicted class). “GXAttack”, “Rnd. Flipping”, and “Rnd. Rewiring” correspond to our attack method, random flipping, and random rewiring, respectively. $\uparrow^{\prime}$ indicates larger values are preferred while $\downarrow^{\prime}$ means the opposite.
@@
From Table II, we have the following main observations:
@@
Observation 2. The type of motifs (namely ground-truth explanations) appears to have minimal impact on both explanation accuracy and attack performance. This can be obtained by comparing the results on Syn3 (Motif =“House”) with those on Syn4 (Motif =“Circle”).
@@
Answer to RQ1. From Table II, we can see that GXAttack can achieve a $\Delta\text{GEA}$ ranging from $0.088$ to $0.267$ while using a very small perturbation budget (less than 7.1 on all datasets) and keeping the predicted labels nearly unchanged (i.e., $\Delta\text{Label}$ is nearly zero) after perturbation.
@@
VI-C Experiments to Answer RQ3: Attack Transferability
@@
Answer to RQ3. From Table III, we can see that some post-hoc explainers, including PGMExplainer, and IG, suffer from poor explanation accuracy on the original graphs, even giving lower accuracy than the Random Explainer. As a result, investigating the attack transferability on these explainers may not be meaningful. On the contrary, like PGExplainer, explainers such as GradCAM, GNNExplainer, SubgraphX, and GNN-LRP can achieve good performance on most original graphs. For those, the attacks optimized for PGExplainer are also harmful, leading to a dramatic degradation of explanation accuracy. For instance, the explanation accuracy of GNNexplainer decreases from 0.711 to 0.397 on Syn2. This demonstrates the transferability of the attacks generated by GXAttack.
VI-D Sensitivity Analysis
Specifically, we perform analysis in terms of $\Delta$GEA, $\Delta$Prob and used budget when varying the following hyperparameters, respectively: 1) the loss trade-off hyperparameter; 2) the maximally allowed perturbation budget (Max Budget) in Figure 2; and 3) the maximal training epochs of attacker in Figure 3.
@@ VI-D Sensitivity Analysis
TABLE III: Transferability of attacks on synthetic datasets Syn1-6 with a typical run (Results on Syn7 are omitted due to excessive computation time). We report the explanation accuracy (i.e., GEA) on the original graph $\rightarrow$ explanation accuracy on the perturbed graph.
Explainer Syn1 Syn2 Syn3 Syn4 Syn5 Syn6
PGExplainer 0.728 $\rightarrow$ 0.510 0.648 $\rightarrow$ 0.347 0.485 $\rightarrow$ 0.304 0.471 $\rightarrow$ 0.311 0.450 $\rightarrow$ 0.285 0.332 $\rightarrow$ 0.225
GradCAM 0.745 $\rightarrow$ 0.551 0.708 $\rightarrow$ 0.397 0.510 $\rightarrow$ 0.300 0.511 $\rightarrow$ 0.310 0.455 $\rightarrow$ 0.278 0.414 $\rightarrow$ 0.241
IG 0.118 $\rightarrow$ 0.104 0.134 $\rightarrow$ 0.103 0.136 $\rightarrow$ 0.101 0.109 $\rightarrow$ 0.092 0.116 $\rightarrow$ 0.085 0.078 $\rightarrow$ 0.065
GNNExplainer 0.738 $\rightarrow$ 0.550 0.711 $\rightarrow$ 0.397 0.539 $\rightarrow$ 0.341 0.514 $\rightarrow$ 0.327 0.483 $\rightarrow$ 0.283 0.431 $\rightarrow$ 0.205
SubgraphX 0.113 $\rightarrow$ 0.015 0.532 $\rightarrow$ 0.375 0.296 $\rightarrow$ 0.141 0.417 $\rightarrow$ 0.282 0.394 $\rightarrow$ 0.312 0.408 $\rightarrow$ 0.305
PGMExplainer 0.015 $\rightarrow$ 0.064 0.046 $\rightarrow$ 0.039 0.031 $\rightarrow$ 0.034 0.036 $\rightarrow$ 0.041 0.038 $\rightarrow$ 0.043 0.037 $\rightarrow$ 0.040
GNN-LRP 0.752 $\rightarrow$ 0.573 0.713 $\rightarrow$ 0.394 0.509 $\rightarrow$ 0.300 0.515 $\rightarrow$ 0.311 0.457 $\rightarrow$ 0.284 0.410 $\rightarrow$ 0.241
RandomExplainer 0.184 $\rightarrow$ 0.146 0.170 $\rightarrow$ 0.096 0.135 $\rightarrow$ 0.102 0.124 $\rightarrow$ 0.090 0.128 $\rightarrow$ 0.100 0.133 $\rightarrow$ 0.086
Figure 2: Sensitivity analysis w.r.t. maximally allowed perturbation budget (Max Budget) on Syn2 using GXAttack.
Attack Budget vs Attack Performance. From Figure 2, we can see that as the Max Budget increases from 1 to 10, $\Delta$GEA and $\Delta$Prob rise sharply, showing large gains from higher allowed budgets. Beyond a (maximally allowed) budget of 10, $\Delta$GEA and $\Delta$Prob both stabilize, indicating no further gains. The underlying reason is that the used budget stop increasing after 10 (maximally allowed) when we set the training epochs as 100. In Tables II and III, we report the results with a maximum budget equal to 15.
Attack Training Epochs vs Attack Performance. From Figure 3, we observe a steady increase in the used budget as epochs increase. However $\Delta$Prob first increases and then decreases (due to overfitting) when increasing the training epochs. Similarly, $\Delta$GEA first increases and then decreases (due to overfitting) with the increase of training epochs. For computation time reason, we only report the results with 100 epochs in Tables II and III. However, it is important to note that the attack performance (in terms of $\Delta$GEA) can be largely improved if we increase the number of training epochs. For instance, $\Delta$GEA can be improved from 0.269 to 0.364 if we increase the training epochs from 100 to 300.
Figure 3: Sensitivity analysis w.r.t. maximal training epochs on Syn2 using GXAttack.
VI-E Property Analysis
We also conducted some property analysis of GXAttack in terms of the prediction confidence, and node degree.
Property Analysis on Prediction Confidence. From Figure 4(a), we can clearly see that nodes in ‘G9’ tend to have lower original explanation accuracy. As a result, as shown in Figure 4(b), the attack performance of GXAttack on these nodes (in ‘G9’) tend to be less successful (i.e., smaller $\Delta\text{GEA}$). This is because the nodes in ‘G9’ leave no much room for perturbations under the constraint that the predictions remain unchanged, namely much smaller probability changes as shown in Figure 4(c).
Property Analysis on Node Degree. From Figure 5(a), we can see that nodes with higher degrees tend to have lower original explanation accuracy. Figure 5(b) shows that the attack performance of GXAttack on these nodes with higher degrees tends to be less successful (i.e., smaller $\Delta\text{GEA}$). Possible reasons are that 1) the original explanation accuracy tends to be lower; and 2) small perturbations on nodes with high degrees have limited impact on the predictions, resulting in smaller probability changes as shown in Figure 5(c).
(a) Prediction confidence vs. original explanation accuracy on Syn2.
(b) Prediction confidence vs. attack performance ($\Delta\text{GEA}$) on Syn2.
(c) Prediction confidence vs. probability change after attacks on Syn2.
Figure 4: Property analysis regarding prediction confidence on Syn2. Other datasets show consistent results and are omitted.
(a) Node degree vs. original explanation accuracy on Syn2.
(b) Node degree vs. attack performance ($\Delta\text{GEA}$) on Syn2.
(c) Node degree vs. probability change after attacks on Syn2.
Figure 5: Property analysis regarding node degree on Syn2. Other datasets show consistent results and thus are omitted.
VII Conclusions
@@ Acknowledgements
Zhong Li and Matthijs van Leeuwen: this publication is part of the project Digital Twin with project number P18-03 of the research programme TTW Perspective, which is (partly) financed by the Dutch Research Council (NWO). Simon Geisler: This research was supported by the Helmholtz Association under the joint research school “Munich School for Data Science – MUDS”.
Appendix A: Projected Gradient Descent
@@
After each gradient update, $\tilde{\mathbf{P}}_{t}\in\mathbb{R}^{n\times n}$, although it must lie in $[0,1]^{n\times n}$ for its interpretation as a Bernoulli random variable. Moreover, we must ensure that sampling $\mathbf{P}_{t}\in\{0,1\}^{n\times n}\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{t})$ will yield a discrete perturbation $\hat{\mathbf{A}}=\mathbf{A}\oplus\mathbf{P}_{t}$ obeying the admissible perturbations $\|\mathbf{A}-\hat{\mathbf{A}}\|_{0}<B$ with sufficient likelihood. Following [46, 47], we ensure that $\mathbb{E}_{\mathbf{P}\sim\text{Bernoulli}(\tilde{\mathbf{P}_{t}})}[\|\mathbf{P}\|_{0}]\leq B$ and then obtain the final perturbation form a small set of samples. To guarantee that $\tilde{\mathbf{P}}_{t}$ parametrizes a valid distribution and that we obey the budget in expectation, we employ the projection $\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$.
@@
$\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}})\leq B]}(\tilde{\mathbf{P}})=\arg\min_{\tilde{\mathbf{S}}}\|\tilde{\mathbf{P}}-\tilde{\mathbf{S}}\|_{F}^{2}$ (21)
subject to $\qquad\text{subject to }\;\sum_{i,j}\tilde{S}_{i,j}\leq B~{}\text{and}~{}\tilde{\mathbf{S}}\in[0,1]^{n\times n}$
@@
for clipping to the hyperplane and projecting back towards the simplex using a bisection search. See [46] for details and the derivation.
Appendix B: Presumptions Checking Results
(a) Syn1
(b) Syn2
(c) Syn3
(d) Syn4
(e) Syn5
(f) Syn6
(g) Syn7